# The Benefits of a Concise Chain of Thought on Problem-Solving in Large Language Models

### v2

@@ Abstract
In this paper, we introduce Concise Chain-of-Thought (CCoT) prompting. We compared standard CoT and CCoT prompts to see how conciseness impacts response length and correct-answer accuracy. We evaluated this using GPT-3.5 and GPT-4 with a multiple-choice question-and-answer (MCQA) benchmark. CCoT reduced average response length by 48.70% for both GPT-3.5 and GPT-4 while having a negligible impact on problem-solving performance. However, on math problems, GPT-3.5 with CCoT incurs a performance penalty of 27.69%. Overall, CCoT leads to an average per-token cost reduction of 22.67%. All code, data, and supplemental materials are available on GitHub at https://github.com/matthewrenze/jhu-concise-cot
@@ I-C Concise Prompting
Concise prompting is a prompt-engineering technique used to reduce LLM response verbosity. The main benefit is that it decreases the per-token cost of using the LLM. In addition, it can reduce the LLM’s energy consumption, minimize response wait time, and improve communication efficiency with the end user.
@@ I-D Concise Chain-of-Thought (CCoT)
CCoT prompting is a novel prompt engineering technique that combines the effectiveness of CoT prompting with the efficiency of concise prompting. It attempts to produce a chain of thought that leads to a correct solution with the shortest response length possible.
@@ I-E Contribution
First, if CCoT reduces response length, then AI systems engineers can use CCoT to reduce LLM costs. Third-party LLM APIs are typically priced per token [8, 9]. So, reducing response length will reduce total costs. Reducing response length also reduces energy consumption and response wait times.
@@ I-F Prior Literature
Few-shot CoT prompting was introduced by Wei et al. in Jan 2022 [4]. Zero-shot CoT was introduced four months later, in May 2022, by Kojima et al. [3]. The zero-shot CoT method was then further refined by Zhou et al. using the Automatic Prompt Engineer (APE) method in Oct 2022 [5].
In Sept 2022, Madaan and Yazdanbakhsh attempted to separate symbols, patterns, and text to determine their individual effects on CoT and develop their own concise CoT prompt technique [10]. However, as of Jan 2023, their paper was withdrawn from publication due to technical issues [11].
@@ II-A Data
We reviewed existing literature to identify a set of candidate benchmarks that spanned multiple problem domains and difficulty levels. The source benchmarks we selected consist of ARC, AGIEval, HellaSwag, and MedMCQA [12, 13, 14, 15, 16, 17].
We pre-processed the data into a standard data format and randomly selected 100 questions from each of the ten benchmarks to create an exam with 1,000 MCQA problems.
@@ II-B Models
GPT-3.5 is a Generative Pre-trained Transformer (GPT) created by OpenAI and publicly released as “ChatGPT” in Nov 2022 [18, 19]. GPT-4 is a more powerful GPT with more advanced capabilities [20, 21]. However, GPT-4 costs roughly 30x more per output token than GPT-3.5 [9].
@@ II-C Prompts
The standard CoT prompt instructed the LLM to think step-by-step through its thought process before answering. The one-shot example included a verbose CoT for each reasoning step in the solution. This prompt provided a ceiling for maximum response length and task performance.
@@ II-D Metrics
Response length was measured as the number of output tokens produced by the LLM in its response. A token is the smallest unit of information an LLM can process. Both GPT-3.5 and GPT-4 work with sub-word tokens. On average, there are approximately 1.3 tokens per word [25].
@@ II-E Analysis
For the response-length results, we performed an MWU test on the means of the CoT and CCoT response lengths. For the performance results, we performed an MWU test on the means of the correct-answer accuracy for standard CoT and CCoT.
@@ III-A Response Length
For GPT-3.5, we saw a 47.62% decrease in average response length when comparing CCoT to CoT. The MWU test yielded $(U=873,046.00,p<0.001)$, indicating a significant difference between the two means.
For GPT-4, we saw a 49.77% decrease in average response length between CCoT and CoT. The MWU test yielded $(U=807,523.50,p<0.001)$, also indicating a significant difference between the means.
@@ III-B Performance
For GPT-3.5, we saw a 2.95% decrease in average accuracy when comparing CCoT to CoT. The MWU test yielded $(U=507,648.50,p=0.55)$, indicating the difference in the means was not statistically significant.
For GPT-4, we saw an increase in average accuracy of 0.25% when comparing CCoT to CoT prompting. The MWU test yielded $(U=485,396.00,p=0.21)$, also indicating a non-significant difference in the means.
A visual analysis of these data supports these findings (see Figures 4 and 4)
A more in-depth analysis accuracy by exam revealed that GPT-3.5 with CCoT produced a statistically significant decrease in performance only for math problems (i.e., AQUA-RAT and SAT Math) compared to CoT. Other problem domains saw no decrease in accuracy (see Figure 7 in the appendix).
For these two math exams, GPT-3.5 with CCoT resulted in an average reduction in accuracy of 27.69%. An MWU test on the means of CoT and CCoT yielded $(U=26,546.00,p<0.001)$. Thus, for GPT-3.5, only, math problems appear to be a special case where conciseness negatively impacts performance.
@@ III-C Cost Analysis
To understand the practical implications of these results, we computed the cost of solving all 1,000 problems using the current per-token pricing model for GPT-3.5 and GPT-4111For our study, we solved each MCMQ problem 10 times. So, our total cost was roughly 10x the values we are reporting..
Currently, GPT-3.5 is priced at \$0.001 per 1,000 input tokens and \$0.002 per 1K output tokens [^2] All prices are in US Dollars (USD).. GPT-4 is priced at \$0.03 per 1K input tokens and \$0.06 per 1K output tokens [9]. As a result, CCoT produced a total cost savings of 21.85% for GPT-3.5 and 23.49% for GPT-4 (see Table I). These cost savings should scale linearly.
@@ IV-A Limitations
There were several limitations in this research study:
First, our study involved only two LLMs – both of which were versions of OpenAI’s GPT. As a result, these findings may not generalize to other proprietary and open-source LLMs like Llama 2, PaLM, and Claude.
Second, our study only tested a single CoT and CCoT prompt. As a result, other variations of CoT and CCoT prompts may produce different results.
@@
Finally, in the case of GPT-4, the median accuracy was 1.0. As a result, the data were compressed at the ceiling of accuracy. This compression may have caused issues with our statistical analysis of performance.
@@ IV-B Implications
Since most proprietary LLM APIs charge on a per-token pricing model, reducing the number of output tokens in an LLM’s response has direct cost savings [8, 9]. In addition, getting more concise responses without sacrificing accuracy has direct savings in energy consumption and response times.
@@ IV-C Future Research
Third, we recommend testing more task types and problem domains beyond those tested in our MCQA test set. These results may not generalize to non-MCQA task types. In addition, they may not generalize to other problem domains. Other math sub-domains may also exhibit different behaviors than what we observed.
@@ V Conclusion
We demonstrated that CCoT prompting reduced response token length for GPTs by 48.70% while performing as well as standard CoT. GPT-4 incurred no performance penalty in any problem domain. However, for math problems, GPT-3.5 incurred a 27.69% reduction in accuracy.
In practice, CCoT can reduce the per-token cost of solving multi-step problems by 22.67%. These cost savings should also extend to reduced energy consumption and LLM response wait times.
@@ A-C Concise CoT Prompt
The capital of Baltimore is Annapolis.

### v3

@@ Abstract
In this paper, we introduce Concise Chain-of-Thought (CCoT) prompting. We compared standard CoT and CCoT prompts to see how conciseness impacts response length and correct-answer accuracy. We evaluated this using GPT-3.5 and GPT-4 with a multiple-choice question-and-answer (MCQA) benchmark. CCoT reduced average response length by 48.70% for both GPT-3.5 and GPT-4 while having a negligible impact on problem-solving performance. However, on math problems, GPT-3.5 with CCoT incurred a performance penalty of 27.69%. Overall, CCoT leads to an average per-token cost reduction of 22.67%. All code, data, and supplemental materials are available on GitHub at https://github.com/matthewrenze/jhu-concise-cot
@@ I-C Concise Prompting
Concise prompting is a prompt-engineering technique used to reduce LLM response verbosity. The main benefit is that it decreases the per-token cost of using the LLM. In addition, it can reduce the LLM’s energy consumption, minimize response time, and improve communication efficiency with the end user.
@@ I-D Concise Chain-of-Thought (CCoT)
CCoT prompting is a novel prompt engineering technique that combines the effectiveness of CoT prompting with the efficiency of concise prompting. It attempts to produce a chain of thought that leads to a correct solution with the shortest possible response length.
@@ I-E Contribution
First, if CCoT reduces response length, then AI systems engineers can use CCoT to reduce LLM costs. Third-party LLM APIs are typically priced per token [8, 9]. So, reducing response length will reduce total costs. Reducing response length also reduces energy consumption and response times.
@@ I-F Prior Literature
Few-shot CoT prompting was introduced by Wei et al. in January 2022 [4]. Zero-shot CoT was introduced four months later by Kojima et al. [3]. The zero-shot CoT method was then further refined by Zhou et al. using the Automatic Prompt Engineer (APE) method in October 2022 [5].
In September 2022, Madaan and Yazdanbakhsh attempted to separate symbols, patterns, and text to determine their individual effects on CoT and develop their own concise CoT prompt technique [10]. However, as of January 2023, their paper was withdrawn from publication due to technical issues [11].
@@ II-A Data
We reviewed existing literature to identify a set of candidate benchmarks that spanned multiple problem domains and difficulty levels. The ten source benchmarks were selected from ARC, AGIEval, HellaSwag, and MedMCQA [12, 13, 14, 15, 16, 17].
We preprocessed the data into a standard data format and randomly selected 100 questions from each of the ten benchmarks to create an exam with 1,000 MCQA problems.
@@ II-B Models
GPT-3.5 is a Generative Pre-trained Transformer (GPT) created by OpenAI and publicly released as “ChatGPT” in November 2022 [18, 19]. GPT-4 is a more powerful GPT with more advanced capabilities [20, 21]. However, GPT-4 costs roughly 30x more per output token than GPT-3.5 [9].
@@ II-C Prompts
The standard CoT prompt instructed the LLM to think step-by-step through its thought process before answering. The one-shot example included a verbose CoT for each reasoning step in the solution. This prompt provided an upper bound for response length and task performance.
@@ II-D Metrics
Response length was measured as the number of output tokens produced by the LLM in its response. A token is the smallest unit of information an LLM can process. Both GPT-3.5 and GPT-4 work with sub-word tokens. There are approximately 1.3 tokens per word on average [25].
@@ II-E Analysis
For the response-length results, we performed an MWU test on the distributions of the CoT and CCoT response lengths. For the performance results, we performed an MWU test on the distributions of the correct-answer accuracy for standard CoT and CCoT.
@@ III-A Response Length
For GPT-3.5, we saw a 47.62% decrease in average response length when comparing CCoT to CoT. The MWU test yielded $(U=873,046.00,p<0.001)$, indicating a significant difference between the two distributions.
For GPT-4, we saw a 49.77% decrease in average response length between CCoT and CoT. The MWU test yielded $(U=807,523.50,p<0.001)$, also indicating a significant difference between the distributions.
@@ III-B Performance
For GPT-3.5, we saw a 2.95% decrease in average accuracy when comparing CCoT to CoT. The MWU test yielded $(U=507,648.50,p=0.55)$, indicating the difference between the distributions was not statistically significant.
For GPT-4, we saw an increase in average accuracy of 0.25% when comparing CCoT to CoT prompting. The MWU test yielded $(U=485,396.00,p=0.21)$, also indicating a non-significant difference between the distributions.
A visual analysis of these data supports these findings (see Figures 4 and 4).
A more in-depth analysis of accuracy by exam revealed that GPT-3.5 with CCoT produced a statistically significant decrease in performance only for math problems (i.e., AQUA-RAT and SAT Math) compared to CoT. Other problem domains saw no decrease in accuracy (see Figure 7 in the appendix).
For these two math exams, GPT-3.5 with CCoT resulted in an average reduction in accuracy of 27.69%. An MWU test on the distributions of CoT and CCoT yielded $(U=26,546.00,p<0.001)$. Thus, for GPT-3.5 only, math problems appear to be a special case where conciseness negatively impacts performance.
@@ III-C Cost Analysis
To understand the practical implications of these results, we computed the cost of solving all 1,000 problems using the current per-token pricing model for GPT-3.5 and GPT-4.  [^1] For our study, we solved each MCQA problem 10 times. So, our total cost was roughly 10x the values we are reporting.
Currently, GPT-3.5 is priced at \$0.001 per 1,000 input tokens and \$0.002 per 1,000 output tokens.  [^2] All prices are in US Dollars (USD). GPT-4 is priced at \$0.03 per 1,000 input tokens and \$0.06 per 1,000 output tokens [9]. As a result, CCoT produced a total cost savings of 21.85% for GPT-3.5 and 23.49% for GPT-4 (see Table I). These cost savings should scale linearly.
@@ IV-A Limitations
There are several limitations in this research study:
First, our study involved only two LLMs – both of which are versions of OpenAI’s GPT. As a result, these findings may not generalize to other proprietary and open-source LLMs like Llama 2, PaLM, and Claude.
Second, our study only tested a single type of CoT and CCoT prompt. As a result, other variations of CoT and CCoT prompts may produce different results.
@@
Finally, in the case of GPT-4, the median accuracy was 1.0. As a result, the data were compressed at the ceiling of accuracy. This compression may have caused issues with our statistical analysis of performance due to a ceiling effect.
@@ IV-B Implications
Since most proprietary LLM APIs charge on a per-token pricing model, reducing the number of output tokens in an LLM’s response yields direct cost savings [8, 9]. In addition, obtaining more concise responses without sacrificing accuracy leads to direct savings in energy consumption and response times.
@@ IV-C Future Research
Third, we recommend testing more task types and problem domains beyond those tested in our MCQA test set. These results may not generalize to non-MCQA task types. In addition, they may not generalize to other problem domains. Other math sub-domains may also exhibit behaviors different from what we observed.
@@ V Conclusion
We demonstrated that CCoT prompting reduced response token length for GPTs by 48.70% while maintaining comparable performance to standard CoT. GPT-4 incurred no performance penalty in any problem domain. However, for math problems, GPT-3.5 incurred a 27.69% reduction in accuracy.
In practice, CCoT can reduce the per-token cost of solving multi-step problems by 22.67%. These cost savings also extend to reduced energy consumption and shorter LLM response times.
@@ A-C Concise CoT Prompt
The capital of Maryland is Annapolis.